OpenStudy (anonymous):

Solve: 2(x + 8) − (x + 1) = 20

OpenStudy (anonymous):

Use the distributive property 2(x + 8) - (x + 1) = 20 2(x) + 2(8) -x - 1 = 20 Can you solve from here?

OpenStudy (uri):

X =3

OpenStudy (uri):

2(x+8)-x+1=20 2x+16-x+1=20 2x-x+16+1=20 x+17=20 x=20-17 x=3